a store mixes kenyan coffee worth $11 per kilogram and venezuelan coffee worth $13 per kilogram. The mixture is to sell for $12 per kilogram. Find how much of each should be used to make a 252-kilogram mixture.
Found 2 solutions by ewatrrr, mananth:
Answer by ewatrrr(24785)


Hi,
Let x represent amount of $13/kg coffee , then (252-x) is the amount of 11/kg coffee
13x + 11(252kg-x) = 12*252kg
2x = 12*252kg - 11*252kg
2x = 252kg
x = 126 kg venezuelan coffee and 126 kg kenyan coffee
checking our answer
13*126 + 11*126 = 1638 + 1386 = 3024 = 12*252

Answer by mananth(16949)

a store mixes kenyan coffee worth $11 per kilogram and venezuelan coffee worth $13 per kilogram. The mixture is to sell for $12 per kilogram. Find how much of each should be used to make a 252-kilogram mixture.
kenyan coffee worth $11 per kilogram : let this be x kg
venezuelan coffee worth $13 per kilogram. : 252-x kg
mixture is to sell for $12 per kilogram. 252-kilogram mixture.
...
11x + 13(252-x)= 12*252
11x+3276-13x = 3024
-2x=-252
x = 171 kg ----------- kenyan coffee
171 kg venezuela coffee
...
m.ananth@hotmail.ca